1. The perimeter of a rectangle is 32 inches, and the area is 60 square inches. find the length and width of the rectangle.
Answer by user_dude2008(1862)

2L+2W=32 ----> L+W=16
LW=60 ----> L=60/W


L+W=16

60/W+W=16

60+w^2=16w

w^2-16w+60=0


(w-10)(w-6)=0

w-10=0 or w-6=0

w=10 or w=6

Answer: Width = 6 Length = 10